OpenStudy (anonymous):

Solve and Check the solution 5x^2-13x=0

OpenStudy (anonymous):

x=0,2.6

OpenStudy (anonymous):

well how did you get that?

OpenStudy (anonymous):

hhhhhhhh

OpenStudy (anonymous):

easy

OpenStudy (anonymous):

by sloving quadratic equetion

OpenStudy (anonymous):

can you show me how? I mean I dont want the answer I want to see how it was solved.

OpenStudy (anonymous):

ok x(5x-13)=0 then x=0 or 5x-13=0 then x=0,13/5

OpenStudy (anonymous):

okay thank you.

OpenStudy (anonymous):

thank u